OpenStudy (annie96):

Which equation can be used to answer the question below? Jeremiah made 14 of the 32 recipes he downloaded from the Internet. What percent of the recipes did he make? (Points : 1) 14x = 32 x = 14(32) 14 = 32x 14=x/32 Question 2.2. Which equation can be used to answer the question below? Josh marked 12% of the e-mails in his Inbox as Important. If he has 125 e-mails in his Inbox, how many did he mark as Important? (Points : 1) 0.12n = 125 125n = 0.12 n = 0.12(125) n=0.12/125

OpenStudy (anonymous):

(14/32) = 0.4375 = 43.75%

OpenStudy (anonymous):

14 = 32x You divide both sides by 32

OpenStudy (anonymous):

12% of 125 = 125*(0.12) = 15

OpenStudy (anonymous):

so n = 0.12(125)